OpenStudy (anonymous):

Please help! 1) Show that the function y=x^21+x^11+13x does not have local maximum or local minimum. 2) Find the values of a and b if the function y=2x^3+ax^2+bx+36 has local maximum when x=-4 and a local minimum when x=5. Thanks! :)

zepdrix (zepdrix):

Mmmm ok, remember how to find max/min (critical) points?

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

So our first derivative tells us about slope. Setting it equal to 0 gives us zero slopes, critical points. Let's try to do that.

zepdrix (zepdrix):

\[\Large\rm y=x^{21}+x^{11}+13x\]\[\Large\rm y'=?\]

OpenStudy (anonymous):

\[y'=21x ^{20}+11x ^{10}+13\]

zepdrix (zepdrix):

Setting equal to zero gives us,\[\Large\rm 0=21x^{20}+11x^{10}+13\]Which I'm going to rewrite like this,\[\Large\rm 0=21(x^{10})^2+11(x^{10})+13\]That step ok? I applied an exponent rule there.

OpenStudy (anonymous):

Yep

zepdrix (zepdrix):

I'll make the substitution \(\Large\rm u=x^{10}\),\[\Large\rm 0=21u^2+11u+13\]

zepdrix (zepdrix):

If we throw this into the quadratic equation we'll find that the discriminant is less than zero (The part under the root), so no real solutions exists for \(\Large\rm u\). Which further implies that no real solutions exists for \(\Large\rm x\).

zepdrix (zepdrix):

That shows us that no critical points exist for this function. Therefore no max/min points will exist.

zepdrix (zepdrix):

Any of that confusing? Do you remember what the quadratic equation looks like?

OpenStudy (anonymous):

Yes, I think i got it! Thanks for the great explanation.

zepdrix (zepdrix):

Mmmm how bout part 2?

OpenStudy (anonymous):

Yes, please!

zepdrix (zepdrix):

So the process starts out the same way, yah? What do you get for your derivative?

OpenStudy (anonymous):

\[y'= 6x^2+2ax+b\] ???

zepdrix (zepdrix):

\[\Large\rm y=2x^3+ax^2+bx+36\]\[\Large\rm y'=6x^2+2ax+b\]Mmmm ok looks good!

zepdrix (zepdrix):

They gave us the exact locations of max and min points of this function. Since \(\Large\rm x=5\) is a maximum of the function, it is also a critical point of the function, and since it is a critical point, it will result in the first derivative equaling zero.

zepdrix (zepdrix):

So plugging in y'=0 and x=5 gives us,\[\Large\rm 0=6(5)^2+2a(5)+b\]

zepdrix (zepdrix):

Since x=5 is a minimum* blah typo, still works out ok though

OpenStudy (anonymous):

okay..

zepdrix (zepdrix):

This gives us an equation involving a, b, and numbers. We want to do the same thing with the x=-4 information, this will give us a `system of 2 equations and 2 unknowns`.

zepdrix (zepdrix):

So this first equation we can write as:\[\Large\rm -150=10a+b\]

zepdrix (zepdrix):

How bout the other one, can you figure that equation out? :)

OpenStudy (anonymous):

Is it -96=-8a+b ?

zepdrix (zepdrix):

Mmm ok very good!

zepdrix (zepdrix):

So we have a system of 2 equations, 2 unknowns.

zepdrix (zepdrix):

Do you remember how to solve a system like that?

OpenStudy (anonymous):

I think I forgot how to do it...

zepdrix (zepdrix):

We would like to combine these equations in some way that causes one of the variables to disappear.